A patient spent 25 minutes waiting in the reception area before seeing a doctor. The visit with the doctor lasted 5 minutes. What is the value-added percentage of the patient?

20%

17%

80%

83%
Answer #1

Total time \(=\) Non-value added \(+\) Value added \(=25+5=30\).

Value added activity (seeing the Doctor) = 5

Hence, Value added percentage = Value added activity/Total time \(\times 100=5 / 30 \times 100=16.66 \%\) (i.e. \(17 \%)\)